OpenStudy (anonymous):

Differentiation: The two tangents to the curve y=x^2 at the points where y=9 intersect at the point P. Find the coordinates of P.

myininaya (myininaya):

this is what I think: if y=9, then 9=x^2 implies x=3 and x=-3 the slope of the tangent line at x=3 and x=-3 can be found by finding the derivative. If you write both equations of each line, you should get that they intersect the y-axis at the same point which is (0,-9)

OpenStudy (anonymous):

wow, i was confused about what the question was asking for and now it makes complete sense.

OpenStudy (anonymous):

thank you!

myininaya (myininaya):

np

OpenStudy (anonymous):

so i got the equations y=-6x-9 and y=6x-9. how do i solve for the intersecting point?

myininaya (myininaya):

well look at the y-intercepts of each line

OpenStudy (anonymous):

for future problems, how would i find the intercept if the linear equations didn't have the same y-intercept?

myininaya (myininaya):

the y-intercepts are the same so they intersect at (0,-9)

myininaya (myininaya):

set them equal to each other. my computer is going slow

OpenStudy (anonymous):

before you crash, know i'm thankful haha

myininaya (myininaya):

lol i can try to make a problem like this where the y intercepts are not the same

OpenStudy (anonymous):

i get it it's ok :)

myininaya (myininaya):

ok